If a²+b²+c²+d²=1 Then find the maximum value of 16abcd

If a²+b²+c²+d²=1
Then find the maximum value of
16abcd

Grade:10

1 Answers

Arun
6 years ago
Dear Ishan
 
Maximum value occur when all of them are equal.
 
Hence
a²+a²+a²+a² = 1
a² = ¼
a = ½ = b = c = d
 
Hence
16*abcd = 16*(1/2)(1/2)(1/2)(1/2) = 16/16 = 1
 
Regards
Arun (askIITians forum expert)